OpenStudy (anonymous):

solve using substitution? y=-3x+1 and 2x+y=1

OpenStudy (anonymous):

2x+(-3x+1)=1 2x-3x+1=1 -1x=0 x=0 plug this into one of the equations -3(0)+1=y 1=y the coordinate of the intercept is (0,1)